OpenStudy (anonymous):

If A is the set of all numbers a^2+4ab+b^2, prove that A is closed under multiplication. I've tried a few different things, but the algebra always gets out of hand. Any help would be appreciated!

OpenStudy (anonymous):

Let r be real number, I will show that there are two numbers a and b so that \[ a^2 + b^2 + 4 a b= r \] Solve the above equation for a in terms of b and r using the quadratic formula. You get \[ a= -2b-\sqrt{3 b^2+r}\\ a= -2b+\sqrt{3 b^2+r}\\ \] choose b so that \( b^2 + r >0\) and take one of the solutions above for a. This choice of a and b gives that \[ a^2 + b^2 + 4 a b= r \] This proves that A is equal to all real numbers and therefore it closed under multiplication.

OpenStudy (anonymous):

Ah, I left out an important bit - a and b are integers.

OpenStudy (anonymous):

This fundamentally changes the problem :(. Sorry about that.

OpenStudy (anonymous):

Can they be negative?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Here is a solution to your problem

OpenStudy (anonymous):

Let \[ w=(a^2 + b^2 + 4 a b)( c^2 +d^2+4 cd)\\ w=a^2 c^2 + 4 a^2 c d + a^2 d^2 + 4 a b c^2 + 16 a b c d + 4 a b d^2 + b^2 c^2 + 4 b^2 c d + b^2 d^2\\ \] We need to find x and y so that \[ w= x^2 + y^2 + 4 x y \] Use the quadratic formula to find x in terms of w and y.

OpenStudy (anonymous):

You find \[ x=-2 y\pm\sqrt{w+3 y^2} \] If we can choose y so that \( w + 3 y^2\) is a perfect square, then we will be done Take y=b c + a d + 4 b d z=a c + 2 b c + 2 a d + 7 b d then \[ w + 3 y^2= z^2 \] Hence one can take \[ x =-2y \pm z\\ x=-2(b c + a d + 4 b d) \pm z=\\-2(b c + a d + 4 b d)\pm ( a c + 2 b c + 2 a d + 7 b d)\\ x= ac - bd \\or\\ x=-a c - 4 b c - 4 a d - 15 b d \] For both cases, you can verify that \[ x^2 + y^2 + 4 x y =\\a^2 c^2 + 4 a b c^2 + b^2 c^2 + 4 a^2 c d + 16 a b c d + 4 b^2 c d + a^2 d^2 + 4 a b d^2 + b^2 d^2 \] This shows that your set is closed